Prev Next

SW&IT Meeting Minutes 2021-10-25

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

  • staff PA due
  • board meeting Wed/Thu

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
  • This week
    • ATT/SW one-slide recap for the board
    • git/github "policy" (= philosophy slides) for new staff
    • close swg recruitment
    • Performance Appraisals
    • msgd/rtdb fedora/python3 for DMS, WIP
    • lbti-sx-ao showtime

YZ

  • Previous:
    • Released TMS automatic data transfer service.
    • Usual TMS meeting, discussions, and software work.
    • LBC alerts list discussion (but not conclusive).
    • Completed performance evaluation form.
  • Next
    • Hopefully resolve the LBC alerts list discussions.
    • Hopefully finish the MODS WCS program verification.
    • Meeting about RM507 machine setup for observers.
    • Misc. LBC and MODS work.
    • Regular TMS meeting and work.
    • Possible TCS/Argos swingarm work pending on telescope time.

BM

  • Last week
    • Installed new ao-supervisor code on LBTI
    • LBTI pause/resume testing
    • SOUL gain ramp / savestatus testing
    • Installed pause/resume version without gain ramp/savestatus change as default on sx-lbtiwfs and sxadsec.
  • This week
    • Work with Allesandro and/or Xianyu Wednesday to update spare BCU2k IP.
    • Merge pause/resume, gain ramp, savestatus changes.
    • Finish up "stage 2" gain ramp changes (delay time, clipping factor and width configurable)
    • Figure out a cleaner deployment strategy (and perhaps install procedure) for LBTI/SOUL.
    • Housekeeper telemetry logging during SBCU timeout.

DE

  • Shopping for storage, proposal submitted. ETA on go-ahead?
  • Windows install on backup server--somehow the image labeled 20H1 on the software site isn't? Or maybe my slipstream attempt broke the image. Will try with fresh installer
  • Windows Server progress, will follow up. Hopefully we can get this this week
  • Disk swaps
  • Safety training
  • New Zoom room for LF/SHARKs got a thumbs up this morning
  • VPN for Fabio thumbs up
  • Will check on JP and see how the Sites transition work is going
  • 507 meeting this afternoon
  • Evaluation

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Met regarding LUCI sub-frame and AO requirements.
    • Lots of discussion of LUCI FITS headers, stale DD values, and AO.
      • The conclusion was that the AO system needs to be responsible for keeping its DD values up to date.
      • LUCI images were checked, and the data in the headers (calibration vs. non-calibration) seems to be acceptable.
    • Found that configuring LUCI in a VM which uses mDNS is tricky; some places need the FQDN (luci-sim.local) while others need the hostname (luci-sim).
    • Fixed the LUCI configuration in my VM and isola-luci; the configuration on osa-luci still needs work.
    • X2Go /XQuartz has been giving me a lot of trouble; it frequently locks up the session, disconnects, and has to be reconnected when running the LUCi GUIs.
      • ssh -Y, however, is unbearably slow, and there doesn't seem to be a better solution to the problem.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Preparing a new LUCI version for testing and deployment that will hopefully fix many of the open RTD issues (which are really Aladin issues).
Topic revision: r10 - 25 Oct 2021, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback